OpenStudy (anonymous):

Can you factor x^2 + 2x ?

OpenStudy (unklerhaukus):

factors of \(x\) are \(x\) and \(x\) factors of \(2x\) are \(2\) and \(x\) pull the common factor out of the brackets

OpenStudy (anonymous):

I don't understand what you mean. I am trying to figure out the rational expression \[\frac{ x^2 + 2x }{ 12x + 54 } - \frac{ 3 - x }{ 8x + 36 }\] So I must start by factoring each equation correct?

OpenStudy (unklerhaukus):

\[x^2+2x=xx+2x=x(x+2)\]

OpenStudy (unklerhaukus):

factors of \(x^2\)***

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):